OpenStudy (anonymous):

what is the solution for -5x-3y=5 and 12y=-20x-20

OpenStudy (radar):

Don't be fooled with this. Note that the two equations can be expressed as: -5x-3y=5 -20x-12y=20 Further note that the second equations is the first one multiplied by 5...What does that tell you about solving ?

OpenStudy (radar):

Change *5 to 4 above.......lol